In a microsoft word document please write an algorithm for the following JAVA code:

import java.util.Scanner;

public class SalesData {
   public static void main(String[] args) {
       int sales[] = new int[12];
       int max = -1, index = 0;
       Scanner sc = new Scanner(System.in);
       int q1 = 0, q2 = 0, q3 = 0, q4 = 0, total = 0;
       for (int i = 0; i < 12; i++) {
           System.out.println("Please enter the data for month " + (i + 1));
           sales[i] = sc.nextInt();
           if (max < sales[i]) {
               max = sales[i];
               index = i;
           }
       }
       for (int i = 0; i < 12; i++) {
           if (i < 3) {
               q1 += sales[i];
           } else if (i < 6) {
               q2 += sales[i];
           } else if (i < 9) {
               q3 += sales[i];
           } else if (i < 12) {
               q4 += sales[i];
           }
           total += sales[i];
       }
       System.out.println("The first quarter total is " + q1);
       System.out.println("The second quarter total is " + q2);
       System.out.println("The third quarter total is " + q3);
       System.out.println("The fourth quarter total is " + q4);
       System.out.println("The Annual sales total is " + total);
       System.out.println("Month " + (index + 1) + " has the highest sales " + max);
   }
}
Answer #1

DECLARE ARR[12],Q1,Q2,Q3,Q4,MAX=-1
FOR I 1 TO 12
   PRINT "ENTER DATA FOR MONTH ",I
   READ ARR[I]
   IF MAX< ARR[I]
       MAX=ARR[I]

FOR I 1 TO 12
   IF I <3
       Q1=Q1+ARR[I]
   IF I <6
       Q2=Q1+ARR[I]
   IF I <9
       Q3=Q1+ARR[I]
   IF I <12
       Q4=Q1+ARR[I]
   TOTAL=TOTAL+ARR[I];
PRINT "Q1 SALES",Q1
PRINT "Q2 SALES",Q2
PRINT "Q3 SALES",Q3
PRINT "Q4 SALES",Q4
PRINT "TOTAL SALES",TOTAL